OpenStudy (anonymous):

~Plz Help~ What is k-2 When k= -4? ___ 12/k My answer is 2...o-o

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

-6

OpenStudy (camerondoherty):

Is this your question? \[\frac{ k-2 }{ \frac{ 12 }{ k } }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait no O_O hold on :3

OpenStudy (anonymous):

\[ \frac{k-2}{\frac{12}{k}}=\frac{-6}{-3}=2 \]

OpenStudy (anonymous):

\[ \frac{\frac{k-2}{12}}{k} = \frac{-1/2}{-4}=\frac{1}{8} \]

OpenStudy (anonymous):

6/3 which is equal to 2

OpenStudy (anonymous):

ty :)

OpenStudy (anonymous):

-6 is the answer